the perimeter of a rectangle is 48". The length is 3" less than twice its width. What are the dimensions of this rectangle
Answer by checkley79(3341)

P=2L+2W
L=2W-3
48=2(2W-3)+2W
48=4W-6+2W
48=6W-6
6W=48+6
6W=54
W=54/6
W=9" ANS. FOR THE WIDTH.
L=2*9-3
L=18-3=15" ANS. FOR THE LENGTH.
PROOF:
48=2*15+2*9
48=30+18
48=48